OpenStudy (anonymous):

Find a ⋅ b. a = 4i - 4j, b = 4i + 5j

OpenStudy (anonymous):

a.b means scalar product (4i-4j).(4i+5j)=16i.i-20j.j(all i.j or j.i are zero)=-4(i.i=j.j=1)